Determine the equation of the line by using the point-slope form. Put the final answer in y = mx + b form.
m = 3 and point (4, −7)

Answer by Boreal(15235)

formula is y-y1=m(x-x1), m slope and (x1, y1) point. Watch signs
y+7=3(x-4)
y=3x-19